OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (7, 0) and a directrix at x = -7.

OpenStudy (anonymous):

y = 1/28 x2 x = 1/28 y2 -28y = x2 y2 = 14x

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

same @Hero

OpenStudy (anonymous):

I'll give a medal even though this isn't my question